OpenStudy (anonymous):

What is the graph of 2x-y=6 ?

OpenStudy (sidsiddhartha):

try to write it in a form like \[\frac{ x }{ a }+\frac{ y }{ b}=1\]

OpenStudy (anonymous):

x-y=3

OpenStudy (sidsiddhartha):

it will look like \[\frac{ x }{ 3 }+\frac{ y }{- 6 }=1\]

OpenStudy (sidsiddhartha):

so x coordinare is=(3,0) and y coordinate is=(0,-6) now just plot it